OpenStudy (dominirican1013):

Please help! Solve tan(x)(tan(x)-1)=0

OpenStudy (anonymous):

\[\Large x = \pm n(\pi) \ , \ x = \frac{\pi}{4}\pm n(\pi) \]

OpenStudy (dominirican1013):

Thank You so much!

OpenStudy (anonymous):

tan(x)(tan(x) - 1) = 0 tan(x) = 0 or tan(x) - 1 = 0 tan(x) = 0 or tan(x) = 1 Solve tan(x) = 0 to get x = 0. Since we can add on multiples of pi, this means that we get x = ±n(pi) where n is any integer Now solve tan(x) = 1 to get x = pi/4. Now add on multiples of pi to get x=pi/4±n(pi) Note: in both cases, you're either using a calculator or the unit circle Now combine the two solutions to get the final answer\[\Large x = \pm n(\pi) \ , \ x = \frac{\pi}{4}\pm n(\pi) \]

OpenStudy (anonymous):

np

OpenStudy (dominirican1013):

Solve on the interval \[[0,2\pi):\] \[2\cos ^{2}x+3cosx+1=0\]